OpenStudy (anonymous):

ʃ x-3/(x+6)^3 dx how do u integrate this ecuation by repeated linear factor

OpenStudy (anonymous):

(9-2x)/(2*(x-6)^2) use integration by parts

OpenStudy (anonymous):

I have a better way.\[\frac{x-3+3-3}{(x-6)^3} = (x-6)^2 + \frac 3{(x-6)^3}\] Now integrate it.

OpenStudy (anonymous):

\[\int\limits \frac{x - 3}{(x + 6)^3}dx\]\[u = x + 6\]\[x = u - 6\]\[dx = du\]\[\int\limits \frac{u - 9}{u^{3}}du= \int\limits (u^{-2} - 9u^{-3}) du\]

OpenStudy (anonymous):

@Ishaan94 where did the +3-3 come from in the numerator?

OpenStudy (anonymous):

curious because i'll be in this section soon

OpenStudy (anonymous):

He just add 0 (+3 - 3), because them we can bundle it like this: \[ \frac{x-6 + 3}{(x-6)^3} = \frac{x-6}{(x-6)^3} + \frac{3}{(x-6)^3}\]that yields what he got. Although, I think it should be a -2, rather than a 2.

OpenStudy (anonymous):

ah i see thank you for the explanation@bmp

OpenStudy (anonymous):

No problem.